# ORACLE-Grasp: Zero-Shot Task-Oriented Robotic Grasping using Large Multimodal Models

### v1

@@ II-B Grasp Candidate Generation
Early stopping criterion. The above process is nominally set to run for $K$ iterations. To reduce processing time and to mitigate the potential effect of multiple grasping regions, we incorporate an early stopping mechanism which also enforces focus on a specific part of the object. This mechanism assesses the spatial distribution of the rectangular mask centers as an indicator of convergence. After $m<K$ iterations, the centroid $\bar{\mathbf{p}}$ of the current set of center points ${\mathbf{p}_{1},\ldots,\mathbf{p}_{m}}$ is computed. Let $10<\rho<$ 1 be a predefined factor, and let $D$ denote the image diagonal. If the maximum distance from any center point to the centroid satisfies
@@
an additional $m$ iterations are performed on a cropped region of the image centered around the object. The cropping bounds are defined based on the largest diagonal that fits within the union of the $m$ selected grid cells ${B_{1},\ldots,B_{m}}$. If condition (2) is again satisfied after these additional iterations, the process terminates early. Otherwise, the procedure continues until all $K$ iterations are completed. This early stopping strategy not only accelerates the overall pipeline execution but also enhances robustness by ensuring that only consistent and reliable grasp candidates contribute to the final decision.
Orientation refinement. When the object’s graspable region is arbitrarily rotated relative to the image axes, the grid structure becomes misaligned, resulting in inadequate coverage and suboptimal mask fitting. This misalignment can significantly affect the accuracy of the process, particularly when the object spans a large portion of the image. To address this, we utilize PCA on the set of center points to estimate their dominant orientation. If the resulting orientation angle $\alpha$ exceeds a predefined threshold $\alpha_{min}$, the image is rotated by $-\alpha$ to better align the object with the grid structure. A new set of $m$ iterations is then executed. The resulting pixel masks are subsequently transformed back into the original image coordinate system via inverse rotation by $\alpha$. This refinement improves the estimation of both the grasp point and its orientation, enhancing the system’s precision and robustness in challenging edge cases.
II-C Grasp Position Refinement
The above process considers approximation of the required grasp point using solely RGB planar perception. Also, it uses several grid approximations that cannot be thoroughly refined due to a long computational time. Hence, the outputted grasp point may slightly deviate from the object pointing at an empty space rather than the solid structure of the object. For example, when the grasp pipeline predicts a position near the center of a kettle’s handle, which is often concave and semi-circular as illustrated in Figure 3, the depth sensor may incorrectly report an exaggerated depth value due to capturing the background through the hollow region.
@@ II-C Grasp Position Refinement
Figure 3: Illustration of a case where the outputted grasp position (circular red marker) $\bar{\mathbf{p}}$ slightly missed the kettle’s handle in the RGB image, leading to wrong depth estimation $z_{\bar{\mathbf{p}}}$. Circular regions (such as $S_{a}$ and $S_{c}$ with radii $r_{a}$ and $r_{c}$, respectively) along the line connecting the camera and original $\bar{\mathbf{p}}$ are checked for improved depth estimation toward refined grasp position $\bar{\mathbf{p}}^{*}$ (circular green marker) at depth $z^{*}$.
@@
III Experiments
In this section, we evaluate the performance of O-Grasp in predicting human-like grasps, and enabling robotic pick up. We adopted LLaMa 3.2-Vision (MLLaMA) [33] as the chosen LMM, with 11 billion parameters for advanced image reasoning tasks. It provides sufficient semantic reasoning power, built upon MLLaMA’s strong vision-language foundation, and its compact size allows for real-time inference on consumer-grade GPUs. Specifically, the model was deployed on a GPU cloud instance equipped with an NVIDIA H200 SXM GPU, 24 CPUs, and 377 GB of RAM. Though cloud inference is faster, preliminary tests show O-Grasp is also deployable on local hardware, with runtime as the main trade-off. Following extensive hyperparameter tuning, we set the key parameters as follows: number of GRP iterations $K=6$, IoU threshold $\gamma=0.4$, early stopping window $m=3$, and stopping factor $\rho=0.3$.
TABLE II: Position and orientation errors for O-Grasp over test-set I
@@ III Experiments
NRMSE (×$\times 10^{2}$) MAE (${}^{\circ})$
@@
O-Grasp w/o SCP 6.8 $\pm$ 6.1 24.3 $\pm$ 20.2
O-Grasp w/ SCP and GRP 5.1 $\pm$ 4.2 19.0 $\pm$ 14.8
Figure 4: Predicted grasps on three RGB images from test-set I, showing position (blue dot) and orientation (blue line). Objects include, from left to right: a vacuum cleaner, an AI-generated drill and a pencil. Red cells indicate regions selected by the LMM as potential grasping areas.
@@
III-A Test-sets
@@
III-B Evaluation metric
The proposed O-Grasp framework outputs a grasp driven by LMM reasoning. Hence, the evaluation of the predicted grasps should be compared to human reasoning, which is task-oriented. For this purpose, each image in the test-sets is annotated by an uninvolved human participant. These annotations were collected using a dedicated Graphical User Interface, allowing annotators to click on potential grasp locations and draw lines representing the intended grasp direction. The participant was instructed to subjectively annotate potential grasp centers along with the corresponding orientations, reflecting natural hand alignments for each grasp. Multiple annotations were allowed to capture different task-dependent grasp strategies - for example, a drill could be grasped by its handle for a drilling task, or by its cylindrical body when intended for handover. This process yields human annotated grasp positions $\{\bar{\mathbf{p}}_{an,1},\bar{\mathbf{p}}_{an,2},\ldots\}$ (in pixel units) and orientations $\{\theta_{an,1},\theta_{an,2},\ldots\}$ for each image in the test-sets, denoting human-like and task-oriented grasps.
The predicted grasps by O-Grasp are evaluated against human annotations. In test-set I, where no metrical data is available and object scales vary, we compute the Normalized Root Mean Squared Error (NRMSE) of the grasp position by dividing the planar distance between the predicted position $\bar{\mathbf{p}}$ and closest annotated position $\bar{\mathbf{p}}_{an,j}$ by the diameter $d$ of the object’s bounding circle: $e=\frac{|\bar{\mathbf{p}}-\bar{\mathbf{p}}_{an,j}|}{d}$. This yields a unit-less, scale-invariant measure of positional accuracy. For test-set II, which includes depth data, we report standard RMSE in metric units. In both test sets, orientation accuracy is evaluated using the Mean Absolute Error (MAE) between the predicted angle $\theta$ and annotated angle $\theta_{an,j}$ across all images.
Figure 6: Position NRMSE and computation runtime with regard to the number of GRP queries $K$ exerted in O-Grasp.
TABLE III: Grasp prediction accuracy and pick up success rates for test-set II
Object Position RMSE (mm) Orientation Success
w/o GR w/ GR MAE () Rate (%)
Coffee bag 34.8 8.9 1.01 80
Phone charger 19.91 4.0 1.51 100
USB cable 21.49 25.0 6.29 100
Brush 3.3 3.3 0.72 100
Cardboard box 29.41 8.1 3.74 80
Briefcase 30.6 21.0 8.22 60
Glue gun 35.11 7.2 15.38 80
Saw 17.5 12.3 2.6 80
Kneading knife 3.0 3.0 7.84 100
Toothpaste 8,9 8.9 5.23 100
Strengthener 15.0 18.8 18.85 80
Dustpan 30.6 11.7 7.37 80
Rubber duck 22.51 5.2 2.64 100
Stapler 10.82 9.2 13.54 100
Plastic cylinder 29.77 6.5 13.05 80
Plastic cup 27.58 16.9 10.72 100
Desk C-Clamp 18.97 19.0 4.76 80
Remote controller 19.21 2.3 4.92 100
Soldering holder 21.14 10.4 7.2 80
Bubble level tool 19.12 4.3 9.37 80
Total 20.9$\pm$9.2 10.3$\pm$6.4 7.5$\pm$5.3 88
* GR - Grasp Refinement based on depth perception (Sec. II-C)
@@ III-B Evaluation metric
Figure 7: Real-world objects from test-set II being grasped by a robot arm using predictions from O-Grasp. For each object, the figure shows a snapshot of the robot executing the grasp and the corresponding RGB image captured by the on-board camera. The demonstrated objects include: (a) cardboard box, (b) coffee bag, (c) six-arm soldering holder, (d) toothpaste tube, (e) dustpan, and (f) Plastic cylinder. In some cases, the initial predicted grasp (in green) was refined using depth data (final grasp shown in red), as described in Section II-C. The green circles mark the clearance region with radius $r(z_{\bar{\mathbf{p}}})$.
III-C Grasp error evaluation
@@
We begin by evaluating grasp accuracy against human annotations in test set I. We include a baseline where the LMM is prompted to return a bounding box coordinates, with the grasp position taken as the box’s center. Though lacking orientation, this serves as a reference for NRMSE comparison. We also compare performance with and without the SCP. Without it, the model receives only the GRP, which lacks contextual cues. Table II reports the NRMSE and MAE for test set I. The baseline results indicate that directly prompting the LMM for a bounding box yields imprecise and often irrelevant regions. Note that grasp refinement was not applied to test-set I due to the absence of depth data, and the reported results exclude this step. The results demonstrate improved accuracy when the SCP is used alongside the GRP, compared to using the GRP alone. This highlights the significant impact of the SCP: by prompting the model to reason about the scene, it encourages context-aware, task-relevant grasp region selection. Figure 4 demonstrates predicted grasps over three RGB images along with an heatmap representing the LMM’s grid selections. Similarly, Figure 5 presents the grasp prediction by O-Grasp along with the human-annotated grasps. The predicted grasps reflect human-like reasoning, focusing on task-relevant regions such as handles.
@@
We further evaluate the inference time required for generating a grasp prediction. First, the duration of a single GRP query is measured as a function of grid size $u\times v$, executing 500 GRP queries over test-set I with $u$ and $v$ sampled uniformly. The average inference time is $0.93\pm 0.07$ seconds per query. The low standard deviation suggests that GRP inference is largely insensitive to grid resolution, allowing flexible adjustment with minimal computational cost. The average total inference time for O-Grasp to predict a complete grasp over the entire test set is $7.73\pm 2.05$ seconds. This latency can be further reduced with faster GPUs or using multiple ones. Notably, 89% of O-Grasp runs terminated early due to the early stopping criterion, underscoring its practical efficiency. To illustrate the trade-off between grasp accuracy and runtime, Figure 6 shows the position NRMSE and runtime on test-set I as a function of the number of GRP queries $K$, revealing improved accuracy at the cost of increased computation.
III-D Robot grasping experiments
@@
In the following experiment, we assess the effectiveness of O-Grasp in guiding a robotic arm to grasp real-world objects. We use a Franka Research 3 robot arm equipped with a simple parallel gripper, and a ZED-Mini camera mounted at the base of the gripper. For each grasp attempt, the arm first positioned the camera to capture an overhead image of the table. O-Grasp then processed this image to predict the grasp point and orientation, which the robot subsequently used to attempt a pick up by simply employing ROS MoveIt motion planning framework.
@@
The evaluation covers all 20 everyday objects from test-set II, with five grasp trials per object. As before, a human annotator labeled the grasp points and orientations without observing the robot’s actual performance. Table III reports the position RMSE (with and without depth-based grasp refinement), orientation MAE, and the robot’s pick up success rate across trials. The availability of depth data in test-set II enables grasp refinement, which significantly improves positional accuracy by allowing precise adjustments before execution. Figure 7 shows snapshots from several grasping trials of the robot, displaying real-time RGB inputs alongside O-Grasp’s predicted grasps.
@@
The results highlight O-Grasp’s robust performance across a diverse set of real-world objects. For instance, the USB cable and toothpaste tube were grasped with high success rates and showed low orientation errors, demonstrating O-Grasp’s effectiveness on both flexible and deformable items. The rubber duck, despite its irregular shape, also achieved high success rate with minimal position error. In contrast, the briefcase resulted in the lowest success rate and a relatively high orientation error. This may be attributed to its large overall size compared to the small handle, potentially requiring additional GRP iterations to identify a more focused grasp region. Similarly, the cardboard box showed the highest position error, likely due to its many potential graspable regions, which increase ambiguity in grasp selection. Despite such challenges, O-Grasp maintains high accuracy and reliability overall, with 88% average success across all grasping attempts.
IV Conclusions

### v2

@@ 3.2 Grasp Candidate Generation
Early stopping criterion. The above process is run for $K$ iterations by default. To reduce processing time and to mitigate the potential effect of multiple grasping regions, we incorporate an early stopping mechanism which also enforces focus on a specific part of the object. This mechanism assesses the spatial distribution of the rectangular mask centers as an indicator of convergence. After $m<K$ iterations, the centroid $\bar{\mathbf{p}}$ of the current set of center points $\mathbf{p}_{1},\ldots,\mathbf{p}_{m}$} is computed. Let $0<\rho<1$ be a predefined factor, and let $D$ denote the image diagonal. If the maximum distance from any center point to the centroid satisfies
@@
an additional $m$ iterations are performed on a cropped region of the image centered around the object. The cropping bounds are defined based on the largest diagonal that fits within the union of the $m$ selected grid cells $B_{1},\ldots,B_{m}$}. If condition (2) is satisfied after these additional iterations, the process terminates early. Otherwise, if the condition is never met, the procedure defaults to selecting the acquired solution after completing all $K$ iterations. If early stopping occurs, $s$ overlapping regions are calculated with existing masks using the IoU method. Consequently, set $\mathcal{B}$ comprises $2m+s$ rectangular masks, which provide the centroids and eigenvectors used to determine grasp position and orientation. This early stopping strategy not only accelerates the overall pipeline execution but also enhances robustness by ensuring that only consistent and reliable grasp candidates contribute to the final decision.
Orientation refinement. Arbitrary rotation of the object’s graspable region relative to the image axes causes grid misalignment, leading to inadequate coverage and suboptimal mask fitting. This misalignment can significantly affect the accuracy of the process, particularly when the object spans a large portion of the image. To address this, we utilize PCA on the set of center points to estimate their dominant orientation. If the resulting orientation angle $\alpha$ exceeds a predefined threshold $\alpha_{min}$, the image is rotated by $\alpha$ to better align the object with the grid structure. A new set of $m$ iterations is then executed. The resulting pixel masks are subsequently transformed back into the original image coordinate system via inverse rotation by $\alpha$. This refinement improves the estimation of both the grasp point and its orientation, enhancing the system’s precision and robustness in challenging edge cases.
The precision of the orientation estimation $\theta$ is derived from the spatial distribution of the selected grid centers. For elongated objects, the set of selected cells $\mathcal{B}_{j}$ forms a point cloud where the variance along the object’s principal axis significantly outweighs the transverse variance. This results in a dominant principal eigenvector, ensuring directional precision even with coarse discretization. To handle non-linear or curved geometries, the iterative refinement process crops and rescales the image around the region of interest. This zooming mechanism effectively performs a local linearization of the object’s geometry, allowing the PCA to compute an accurate grasp axis on a localized, nearly-linear segment of the object. This combination of global reasoning and local geometric analysis allows ORACLE-Grasp to maintain high directional accuracy across diverse object profiles.
3.3 Grasp Position Refinement
The above process considers approximation of the required grasp point using RGB planar perception alone. Also, it uses several grid approximations that cannot be thoroughly refined due to a long computational time. Hence, the resulting grasp point may slightly deviate from the object, pointing instead to empty space rather than the solid structure of the object. For example, when the grasp pipeline predicts a position near the center of a kettle’s handle, which is often concave and semi-circular as illustrated in Figure 3, the depth sensor may incorrectly report an exaggerated depth value due to capturing the background through the hollow region.
@@ 3.3 Grasp Position Refinement
Figure 3: Illustration of a case where the predicted grasp position (circular red marker) $\bar{\mathbf{p}}$ slightly missed the kettle’s handle in the RGB image, leading to wrong depth estimation $z_{\bar{\mathbf{p}}}$. Circular regions (such as $S_{a}$ and $S_{c}$ with radii $r_{a}$ and $r_{c}$, respectively) along the line connecting the camera and original $\bar{\mathbf{p}}$ are checked for improved depth estimation toward refined grasp position $\bar{\mathbf{p}}^{*}$ (circular green marker) at depth $z^{*}$.
@@
4 Experiments
In this section, we evaluate the performance of ORACLE-Grasp in predicting human-like grasps, and enabling robotic pick-up. For our experimental validation, we adopt LLaMa 3.2-Vision (MLLaMA) [Grattafiori2024] as the primary LMM. This choice was driven by the objective of utilizing state-of-the-art open-weight models to ensure the framework remains accessible, reproducible, and deployable on local hardware without reliance on proprietary APIs. While larger closed-source models (such as GPT-4o) consistently achieve higher scores on standard multimodal benchmarks, we utilize the 11B-parameter LLaMa model as a reliable baseline. Demonstrating the effectiveness of ORACLE-Grasp with this more compact and open-source backbone suggests that the framework’s performance would likely be further enhanced if integrated with more powerful, high-parameter proprietary models.
The MLLaMA model was deployed on a GPU cloud instance equipped with an NVIDIA H200 SXM GPU, 24 CPUs, and 377 GB of RAM. Though cloud inference is faster, preliminary tests show ORACLE-Grasp is also deployable on local hardware, with runtime as the main trade-off. Following extensive hyperparameter tuning, we set the key parameters as follows: number of GRP iterations $K=6$, IoU threshold $\gamma=0.4$, early stopping window $m=3$, and stopping factor $\rho=0.3$.
Table 2: Position and orientation errors for ORACLE-Grasp over test-set I
@@ 4 Experiments
NRMSE ($\times 10^{2}$) MAE ($^{\circ})$
@@
ORACLE-Grasp w/o SCP 6.8 $\pm$ 6.1 24.3 $\pm$ 20.2
ORACLE-Grasp w/ SCP and GRP 5.1 $\pm$ 4.2 19.0 $\pm$ 14.8
4.1 Test-sets
[…]
4.2 Evaluation metric
The ORACLE-Grasp framework produces grasps guided by the reasoning capabilities of an LMM. Therefore, evaluating the predicted grasps should involve comparing them to human judgments, which typically align with the object’s affordances. For this purpose, each image in the test-sets is annotated by an uninvolved human participant. These annotations were collected using a dedicated Graphical User Interface, allowing annotators to click on potential grasp locations and draw lines representing the intended grasp direction. The participant was instructed to subjectively annotate potential grasp centers along with the corresponding orientations, reflecting natural hand alignments for each grasp. Multiple annotations were allowed to capture different task-dependent grasp strategies - for example, a drill could be grasped by its handle for a drilling task, or by its cylindrical body when intended for handover. This process yields human annotated grasp positions $\{\bar{\mathbf{p}}_{an,1},\bar{\mathbf{p}}_{an,2},\ldots\}$ (in pixel units) and orientations $\{\theta_{an,1},\theta_{an,2},\ldots\}$ for each image in the test-sets, denoting human-like and affordance-aligned grasps.
The predicted grasps by ORACLE-Grasp are evaluated against human annotations. In test-set I, where no metrical data is available and object scales vary, we compute the Normalized Root Mean Squared Error (NRMSE) of the grasp position by dividing the planar distance between the predicted position $\bar{\mathbf{p}}$ and closest annotated position $\bar{\mathbf{p}}_{an,j}$ by the diameter $d$ of the object’s bounding circle: $e=\frac{|\bar{\mathbf{p}}-\bar{\mathbf{p}}_{an,j}|}{d}$. This yields a unit-less, scale-invariant measure of positional accuracy. For test-set II, which includes depth data, we report standard RMSE in metric units. In both test-sets, orientation accuracy is evaluated using the Mean Absolute Error (MAE) between the predicted angle $\theta$ and annotated angle $\theta_{an,j}$ across all images.
Figure 4: Predicted grasps on three RGB images from test-set I, illustrating the predicted grasp position (blue dot) and orientation (blue line). Objects include, from left to right: a vacuum cleaner, an AI-generated drill and a pencil. Red cells indicate regions selected by the LMM as potential grasping areas.
@@ 4.2 Evaluation metric
Figure 6: Position NRMSE and computation runtime with regard to the number of GRP queries $K$ exerted in ORACLE-Grasp.
Table 3: Grasp prediction accuracy and pick-up success rates for test-set II (Ablation Study)
Object Position RMSE (mm) Orientation MAE () Success rate (%)
w/o GR w/ GR
w/o OR w/o BE Full w/o OR Full
Coffee bag 34.8 8.9 9.9 8.9 1.01 1.01 80
Phone charger 19.91 8.32 14.7 4.0 6.50 1.51 100
USB cable 21.49 25.0 10.9 25.0 6.29 6.29 100
Brush 3.3 14.3 21.8 3.3 11.72 0.72 100
Cardboard box 29.41 8.1 17.0 8.1 9.92 3.74 80
Briefcase 30.6 21.0 32.3 21.0 8.22 8.22 60
Glue gun 35.11 7.2 26.6 7.2 15.38 15.38 80
Saw 17.5 16.88 8.6 12.3 19.5 2.6 80
Kneading knife 3.0 10.42 7.8 3.0 7.84 7.84 100
Toothpaste 8.9 8.9 10.5 8.9 5.23 5.23 100
Strengthener 15.0 18.8 16.4 18.8 18.85 18.85 80
Dustpan 30.6 20.68 7.8 11.7 28.92 7.37 80
Rubber duck 22.51 5.2 32.5 5.2 2.64 2.64 100
Stapler 10.82 12.81 13.4 9.2 17.14 13.54 100
Plastic cylinder 29.77 6.5 11.8 6.5 13.05 13.05 80
Plastic cup 27.58 16.9 8.4 16.9 10.72 10.72 100
Desk C-Clamp 18.97 26.18 18.8 19.0 0.15 4.76 80
Remote controller 19.21 2.3 9.4 2.3 4.92 4.92 100
Soldering holder 21.14 20.49 34.7 10.4 7.46 7.2 80
Bubble level tool 19.12 13.95 17.5 4.3 9.58 9.37 80
@@
* GR - Grasp Refinement (Sec. 3.3)
* OR - Orientation Refinement
* BE - Brief Explanation
@@
Figure 7: Real-world objects from test-set II being grasped by a robot arm using predictions from ORACLE-Grasp. For each object, the figure shows a snapshot of the robot executing the grasp and the corresponding RGB image captured by the on-board camera. The demonstrated objects include: (a) cardboard box, (b) coffee bag, (c) six-arm soldering holder, (d) toothpaste tube, (e) dustpan, and (f) Plastic cylinder. In some cases, the initial predicted grasp (in green) was refined using depth data (final grasp shown in red), as described in Section $3.$ The green circles mark the clearance region with radius $r(z_{\bar{\mathbf{p}}})$.
4.3 Grasp error evaluation
@@
We begin by evaluating grasp accuracy against human annotations in test-set I. We include a baseline where the LMM is prompted to return a bounding box coordinates, with the grasp position taken as the box’s center. Though lacking orientation, this serves as a reference for NRMSE comparison. We also compare performance with and without the SCP. Without it, the model receives only the GRP, which lacks contextual cues. Table 2 reports the NRMSE and MAE for test-set I. The baseline results indicate that directly prompting the LMM for a bounding box yields imprecise and often irrelevant regions. Note that grasp refinement was not applied to test-set I due to the absence of depth data, and the reported results exclude this step. The results demonstrate improved accuracy when the SCP is used alongside the GRP, compared to using the GRP alone. This highlights the significant impact of the SCP: by prompting the model to reason about the scene, it encourages context-aware, task-relevant grasp region selection. Figure 4 demonstrates predicted grasps over three RGB images along with a heatmap representing the LMM’s grid selections. Similarly, Figure 5 presents the grasp prediction by ORACLE-Grasp along with the human-annotated grasps. The predicted grasps reflect human-like reasoning, focusing on task-relevant regions such as handles.
@@
We further evaluate the inference time required for generating a grasp prediction. First, the duration of a single GRP query is measured as a function of grid size $u\times v$, executing 500 GRP queries over test-set I with $u$ and $v$ sampled uniformly. The average inference time is $0.93\pm 0.07$ seconds per query. The low standard deviation suggests that GRP inference is largely insensitive to grid resolution, allowing flexible adjustment with minimal computational cost. The average total inference time for ORACLE-Grasp to predict a complete grasp over the entire test-set is $7.73\pm 2.05$ seconds. This latency can be further reduced with faster GPUs or using multiple ones. Notably, 89% of ORACLE-Grasp runs terminated early due to the early stopping criterion, underscoring its practical efficiency. To illustrate the trade-off between grasp accuracy and runtime, Figure 6 shows the position NRMSE and runtime on test-set I as a function of the number of GRP queries $K$, revealing improved accuracy at the cost of increased computation.
4.4 Robot grasping experiments
@@
In the following experiment, we assess the effectiveness of ORACLE-Grasp in guiding a robotic arm to grasp real-world objects. We use a Franka Research 3 robot arm equipped with a simple parallel gripper, and a ZED-Mini camera mounted at the base of the gripper. For each grasp attempt, the arm first positioned the camera to capture an overhead image of the table. ORACLE-Grasp then processed this image to predict the grasp point and orientation, which the robot subsequently used to attempt a pick-up by simply employing ROS MoveIt motion planning framework.
@@
The evaluation covers all 20 everyday objects from test-set II, with five grasp trials per object. As before, a human annotator labeled the grasp points and orientations without observing the robot’s actual performance. Table 3 summarizes the experimental results, detailing the position RMSE, orientation MAE, and the pick-up success rate across all trials. To evaluate the individual contributions of our system components, we present an ablation study comparing the performance of the full framework against configurations without: (i) depth-based Grasp Refinement (GR), (ii) the Brief Explanation (BE) CoT format in the GRP, and (iii) the Orientation Refinement (OR) module. The availability of depth data in test-set II enables grasp refinement, which significantly improves positional accuracy by allowing precise adjustments before execution. Figure 7 shows snapshots from several grasping trials of the robot, displaying real-time RGB inputs alongside ORACLE-Grasp’s predicted grasps.
@@
Experimental results underscore ORACLE-Grasp’s robust performance across a diverse range of real-world objects, highlighting significant accuracy gains from the iterative grasp refinement, orientation refinement and the CoT reasoning facilitated by the ’Brief Explanation’ field in the GRP. By forcing the model to articulate its functional rationale, the framework achieves higher spatial precision in its selections. For instance, the USB cable and toothpaste tube were grasped with high success rates and showed low orientation errors, demonstrating ORACLE-Grasp’s effectiveness on both flexible and deformable items. The rubber duck, despite its irregular shape, also achieved high success rate with minimal position error. In contrast, the briefcase resulted in the lowest success rate and a relatively high orientation error. This may be attributed to its large overall size compared to the small handle, potentially requiring additional GRP iterations to identify a more focused grasp region. Similarly, the cardboard box showed the highest position error, likely due to its many potential graspable regions, which increase ambiguity in grasp selection. Despite such challenges, ORACLE-Grasp maintains high accuracy and reliability overall, with 88% average success across all grasping attempts. Finally, while the Early Stopping mechanism does not influence the final accuracy, it serves as a critical efficiency module. In our experiments, it achieved a mean reduction in inference time of 27.48% (±24.70%), allowing the system to converge on an optimal grasp candidate more rapidly.
The impact of the depth-based grasp refinement module is quantified in the results. By comparing the positional RMSE with and without this refinement, we observe a significant enhancement in accuracy across diverse geometries. For instance, for the ’Coffee bag’ and ’Cardboard box,’ the RMSE was reduced by over 70%. It is important to note that since the grasp refinement module operates on local depth data to refine the 3D contact point, it does not alter the grasp orientation $\theta$ derived during the reasoning phase. Furthermore, the reduction in positional error directly correlates with the high success rates observed; without grasp refinement, the initial 2D predictions occasionally target hollow or concave regions, which would result in physical grasp failure. These results validate grasp refinement as a critical bridge between high-level LMM reasoning and precise physical execution.
5 Conclusions